OpenStudy (anonymous):

Give the exact value of each of the following: 2 sin π/6

Parth (parthkohli):

\({\pi \over 6} = {180 \over 6} = 30^{\circ}\)

Parth (parthkohli):

\(\sin\) is 0.5 at 30 degrees, so \(\sin {\pi \over 6} = {1 \over 2}\). Now we have to find \(2 \sin {\pi \over 6}\), so that becomes \(2 \times {1 \over 2}\).

Parth (parthkohli):

Do you need any more help?

OpenStudy (anonymous):

- cos π/4. I got stuck on this questions